An ice making machine operates on the ideal vapor compression cycle, using R-134a. The refrigerant

enters the compressor as saturated vapor at 140kPa and leaves the condenser as saturated liquid at

700kPa. Water enters the ice machine at 13 0 C and leaves as ice -4 0 C. For an ice production rate of 9kg/h,

determine the power input to the ice machine.


1
Expert's answer
2021-06-30T00:31:02-0400

From the data given

"h_1 = 238.97 kJ\/kg; h_2=267.51kJ\/kg; h_3=h_4=77.67 kJ\/kg"

Cooling load

"\dot{Q_L}=\dot{m_{ice}} \triangle{h_{ice}} = \frac{9}{3600}*455.896 =1.13974 kJ"

Refrigerant mass

"\dot{m_R}=\frac{\dot{Q_L}}{h_1-h_4}=\frac{1.13974}{238.97-77.67}=0.007066 kg/s"

The power input

"W_{in}=\dot{m_R}(h_2-h_1) = 0.007066(267.51-238.97)=0.20166 kJ/s = 0.27043 hp"